# First-Order Sparse Convex Optimization: Better Rates with Sparse Updates

### v1

@@ Abstract
In was recently established that for convex optimization problems with a sparse optimal solution (may it be entry-wise sparsity or matrix rank-wise sparsity) it is possible to have linear convergence rates which depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipchitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of the optimal solution. However, beyond the improved convergence rate, these methods are unable to leverage the sparsity of optimal solutions towards improving also the runtime of each iteration, which may still be prohibitively high for high-dimensional problems. In this work, we establish that linear convergence rates which depend on this improved condition number can be obtained using only sparse updates, which may result in overall significantly improved running times. Moreover, our methods are considerably easier to implement.
@@ 1 Introduction
where $f$ is convex and continuously differentiable, and we assume that Problem (1) indeed admits a unique optimal solution ${\mathbf{x}}^{*}$ such that $\textrm{nnz}({\mathbf{x}}^{*})<<n$, where $\textrm{nnz}(\cdot)$ denotes the number of non-zero entries.
@@
$\forall{\mathbf{x}}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\sum_{i=1}^{n}{\mathbf{x}}_{i}-\sum_{j=1}^{n}{\mathbf{x}}^{*}_{j}}\right)^{2}.$ (4)
@@
The recent works [16, 14]111while [16, 14] consider a stochastic convex optimization setting in which the objective function $f$ is given as an expectation w.r.t. unknown distribution from which we can sample, here our focus is on the standard deterministic first-order convex optimization modelindeed lead to convergence rates of the form $\tilde{O}\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ [16] and even an accelerated rate of the form $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$ [14] (the $\tilde{O}(\cdot)$ notation hides certain poly-logarithmic factors), by applying a restarted non-Euclidean gradient method222these methods rely on Bregman proximal gradient computations with the Bregman distance induced by a squared $\ell_{p}$ norm for an appropriate choice of $p$ that best leverages the strong convexity w.r.t. the $\ell_{1}$ norm or its accelerated variant, together with hard-thresholding steps to sparsity intermediate iterates, see for instance Corollary 5 in [14].
@@
Compared to [16, 14], we do have an additional mild assumption, which in the case of Problems (1) and (2) above, requires that the optimal solution is on the boundary in the sense that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ or $\textrm{Tr}({\mathbf{X}}^{*})=R$, respectively. We consider this to be quite mild since in the context of sparse recovery it is not likely that the optimal solution lies in the interior (in particular it implies that the norm regularization is redundant). Additionally, in case of the matrix Problem (2), we shall have an additional term in the condition number, namely the condition number will be of the form: $\left({s+\frac{R^{2}}{\lambda_{\min}({\mathbf{X}}^{*})^{2}}}\right)\frac{\beta_{1}}{\alpha_{2}}$, where $\lambda_{\min}({\mathbf{X}}^{*})$ denotes the smallest non-zero eigenvalue of the optimal solution.
@@ 3.2 Bounded positive semidefinite cone setup
$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq\left({2\sqrt{\textrm{rank}({\mathbf{X}})}+\frac{2\|{{\mathbf{Y}}}\|_{2}}{\lambda_{\min}({\mathbf{X}})}}\right)\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2},$
where $\lambda_{\min}({\mathbf{X}})>0$ denotes the smallest non-zero eigenvalue of ${\mathbf{X}}$.
Let $s$ be such that $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~{}|~{}\textrm{rank}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~{}|~{}{\mathbf{X}}\succeq 0,~{}\|{{\mathbf{X}}}\|_{1}\leq R,~{}\textrm{rank}({\mathbf{X}})\leq s\}$. In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=\left({2\sqrt{s}+\frac{2R}{\lambda_{\min}({\mathbf{X}}^{*})}}\right)^{2}$ (note that for any feasible ${\mathbf{Y}}$ it holds that $\|{{\mathbf{Y}}}\|_{2}\leq R$).
Here the hard-thresholding step in Algorithm 1 (computing the point $\widehat{{\mathbf{x}}}_{t}$) is done by computing the rank-$s$ approximation of the current iterate ${\mathbf{x}}_{t}$, which requires computing only the top $s$ components in the SVD of ${\mathbf{x}}_{t}$ (and discarding the rest).
@@
Overall, aside from the gradient computation, the runtime of each iteration of Algorithm 1 is dominated by computing two rank-$s$ eigen-decompositions of a $n\times n$ symmetric matrix, which is far more efficient than a full-rank eigen-decomposition once $s<<n$.
@@
When applied to Problem (9) with an optimal solution ${\mathbf{X}}^{*}$ which satisfies $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and $\textrm{Tr}({\mathbf{x}}^{*})=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\left({\sqrt{s}+\frac{R}{\lambda_{\min}({\mathbf{X}}^{*})}}\right)^{2}\frac{\beta_{1}}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside of the gradient computations, the runtime of each iteration is dominated by two rank-$s$ eigen-decompositions of $n\times n$ symmetric matrices.
Comparing the rate in Corollary 2 to the state-of-the-art non-Euclidean accelerated gradient method in [14], which has convergence rate $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$, we can see that our rate is not only unaccelerated (i.e., linear in the condition number $\frac{s\beta_{1}}{\alpha_{2}}$ as opposed to linear in its square-root), but we also have an additional term of $\left({\frac{R}{\lambda_{\min}({\mathbf{X}}^{*})}}\right)^{2}$, which may be significantly larger than 1. On the other-hand, the method in [14] requires a full eigen-decomposition of a $n\times n$ matrix each iteration, which is often prohibitive for large $n$.
@@ 4 Some Numerical Evidence
When implementing our Algorithm 1, we assumed the exact sparsity of ${\mathbf{x}}^{*}$ is known, i.e., we set $s=\textrm{nnz}({\mathbf{x}}^{*})$. Additionally, observing that the theoretical choice of step-size corresponding to Theorem 1, $\eta=\frac{\alpha_{2}}{6\beta_{1}{}s}$ (see Section 3.1) is too conservative in practice, we used a larger step-size $\tilde{\eta}=\frac{\alpha_{2}}{\beta_{1}{}s}$, when computing the point ${\mathbf{v}}_{t}$ on each iteration $t$. Additionally, instead of using again the step-size $\tilde{\eta}$ when taking the convex combination to produce the next iterate ${\mathbf{x}}_{t+1}$, we took a slightly larger step-size $2\tilde{\eta}$ (which is still in $(0,1)$). For the V-FISTA method we observed that even trying to slightly tune the prescribed theoretical step-size makes the method divergent, and so we stuck to the original implementation in [2].
@@ 5 Open Questions
We conclude with several open questions which we believe are of interest. First, it is interesting if our approach could be extended to also handle norm-regulairzed problems instead of norm-constrained problems. Second, it is interesting if our approach could be accelerated, i.e., to yield worst-case number of iterations which scales only with $\sqrt{\frac{s\beta_{1}}{\alpha_{2}}}$ while maintaining the sparse updates. Finally, when comparing the matrix Problem (9) with Problem (8), it is interesting if the additional term of $R^{2}/\lambda_{\min}({\mathbf{X}}^{*})^{2}$ in the condition number could be removed.
@@ Appendix B Proof of Lemma 2
Denote $k=\textrm{rank}({\mathbf{X}})$. Let ${\mathbf{P}}\in\mathbb{S}^{n}$ denote the projection matrix onto the subspace $\textrm{Im}({\mathbf{X}})$ and let ${\mathbf{P}}_{\perp}\in\mathbb{S}^{n}$ denote the projection matrix onto the orthogonal complement of $\textrm{Im}({\mathbf{X}})$. Also, let ${\mathbf{Q}}\in\mathbb{S}^{n}$ denote the projection matrix onto subspace spanned by the top $k$ eigenvectors of ${\mathbf{Y}}$ (corresponding to the $k$ largest signed eigenvalues). Recall that $\|{\cdot}\|_{1}$ and $\|{\cdot}\|_{2}$ denote the trace norm and Euclidean norm in $\mathbb{S}^{n}$, respectively. In the following we also let $\sigma_{i}({\mathbf{Z}})$ denote the $i$th largest singular value of some real matrix ${\mathbf{Z}}$. It holds that,
$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}$ $=\|{({\mathbf{P}}+{\mathbf{P}}_{\perp})({\mathbf{X}}-{\mathbf{Y}})}\|_{1}$
$=\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})+{\mathbf{P}}_{\perp}{\mathbf{Y}}}\|_{1}$
$\leq\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})}\|_{1}+\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}}\|_{1}$
$\leq\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}}\|_{1}$
$=\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}-{\mathbf{P}}{\mathbf{Y}}}\|_{1}$
$\leq\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}-{\mathbf{Q}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$=\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}}\|_{1}-\|{{\mathbf{Q}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\underset{(a)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{X}}}\|_{1}-\|{{\mathbf{Q}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\underset{(b)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{X}}}\|_{1}-\|{{\mathbf{P}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\underset{(c)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\leq\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\sqrt{k}\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})}\|_{2}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\leq 2\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\underset{(d)}{\leq}2\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}}\|_{2}\|{{\mathbf{Q}}-{\mathbf{P}}}\|_{2},$
where (a) follows from the assumption of the lemma, (b) follows since on one hand $\|{{\mathbf{Q}}{\mathbf{Y}}}\|_{1}=\sum_{i=1}^{k}\sigma_{i}({\mathbf{Y}})$, and on the other hand, using von Neumann’s inequality for the singular values (see for instance [19]), we have that $\|{{\mathbf{P}}{\mathbf{Y}}}\|_{1}=\sum_{i=1}^{n}\sigma_{i}({\mathbf{P}}{\mathbf{Y}})\leq\sum_{i=1}^{n}\sigma_{i}({\mathbf{P}})\sigma_{i}({\mathbf{Y}})\leq\sum_{i=1}^{k}\sigma_{i}({\mathbf{Y}})$, where we have used the fact that $\textrm{rank}({\mathbf{P}})\leq k$ and $\|{{\mathbf{P}}}\|_{\infty}=1$. Inequality (c) follows since ${\mathbf{X}}={\mathbf{P}}{\mathbf{X}}$, and (d) follows again from von Neumann’s inequality and the Cauchy-Schwarz inequality:
$\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$ $\leq\sum_{i=1}^{n}\sigma_{i}({\mathbf{Q}}-{\mathbf{P}})\sigma_{i}({\mathbf{Y}})$
$\leq\sqrt{\sum_{i=1}^{n}\sigma_{i}({\mathbf{Q}}-{\mathbf{P}})^{2}}\sqrt{\sum_{i=1}^{n}\sigma_{i}({\mathbf{Y}})^{2}}$
$=\|{{\mathbf{Y}}}\|_{2}\cdot\|{{\mathbf{Q}}-{\mathbf{P}}}\|_{2}.$
Using the Davis-Kahan Theorem (see for instance Lemma A.1 in [11]) we have that,
$\|{{\mathbf{Q}}-{\mathbf{P}}}\|_{2}\leq\frac{2\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}}{\lambda_{k}({\mathbf{X}})}.$
Thus, we conclude that
$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq\left({2\sqrt{k}+\frac{2\|{{\mathbf{Y}}}\|_{2}}{\lambda_{k}({\mathbf{X}})}}\right)\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}.$

### v2

@@ Abstract
It was recently established that for convex optimization problems with a sparse optimal solution (may it be entry-wise sparsity or matrix rank-wise sparsity) it is possible to have linear convergence rates which depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipchitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of the optimal solution. However, beyond the improved convergence rate, these methods are unable to leverage the sparsity of optimal solutions towards improving also the runtime of each iteration, which may still be prohibitively high for high-dimensional problems. In this work, we establish that linear convergence rates which depend on this improved condition number can be obtained using only sparse updates, which may result in overall significantly improved running times. Moreover, our methods are considerably easier to implement.
@@ 1 Introduction
[…]
@@
$\forall{\mathbf{x}}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\sum_{i=1}^{n}{\mathbf{x}}_{i}-\sum_{j=1}^{n}{\mathbf{x}}^{*}_{j}}\right)^{2}.$
@@
The recent works [16, 14]111while [16, 14] consider a stochastic convex optimization setting in which the objective function $f$ is given as an expectation w.r.t. unknown distribution from which we can sample, here our focus is on the standard deterministic first-order convex optimization model indeed lead to convergence rates of the form $\tilde{O}\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ [16] and even an accelerated rate of the form $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$ [14] (the $\tilde{O}(\cdot)$ notation hides certain poly-logarithmic factors), by applying a restarted non-Euclidean gradient method222these methods rely on Bregman proximal gradient computations with the Bregman distance induced by a squared $\ell_{p}$ norm for an appropriate choice of $p$ that best leverages the strong convexity w.r.t. the $\ell_{1}$ norm or its accelerated variant, together with hard-thresholding steps to sparsify intermediate iterates, see for instance Corollary 5 in [14].
@@
Compared to [16, 14], we do have an additional mild assumption, which in the case of Problems (1) and (2) above, requires that the optimal solution is on the boundary in the sense that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ or $\textrm{Tr}({\mathbf{X}}^{*})=R$, respectively. We consider this to be quite mild since in the context of sparse recovery it is not likely that the optimal solution lies in the interior (in particular it implies that the norm regularization is redundant).
@@ 3.2 Bounded positive semidefinite cone setup
$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq 4\sqrt{\textrm{rank}({\mathbf{X}})}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}.$
Let $s$ be such that $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~\textrm{rank}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\|{{\mathbf{X}}}\|_{1}\leq R,~\textrm{rank}({\mathbf{X}})\leq s\}$. In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=16s$.
[…]
@@
[…]
@@
[…]
@@ 4 Some Numerical Evidence
When implementing our Algorithm 1, we assumed the exact sparsity of ${\mathbf{x}}^{*}$ is known, i.e., we set $s=\textrm{nnz}({\mathbf{x}}^{*})$. Additionally, observing that the theoretical choice of step-size corresponding to Theorem 1, $\eta=\frac{\alpha_{2}}{6\beta_{1}{}s}$ (see Section 3.1) is too conservative in practice, we used a larger step-size $\tilde{\eta}=\frac{\alpha_{2}}{\beta_{1}{}s}$, when computing the point ${\mathbf{v}}_{t}$ on each iteration $t$. Additionally, instead of using again the step-size $\tilde{\eta}$ when taking the convex combination to produce the next iterate ${\mathbf{x}}_{t+1}$, we took a slightly larger step-size $2\tilde{\eta}$ (which is still in (0,1)(0,1)( 0 , 1 )). For the V-FISTA method we observed that even trying to slightly tune the prescribed theoretical step-size makes the method divergent, and so we stuck to the original implementation in [2].
@@ 5 Open Questions
We conclude with two open questions which we believe are of interest. First, it is interesting if our approach could be extended to also handle norm-regulairzed problems instead of norm-constrained problems. Second, it is interesting if our approach could be accelerated, i.e., to yield worst-case number of iterations which scales only with $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ while maintaining the sparse updates.
@@ Appendix B Proof of Lemma 2
Denote $k=\textrm{rank}({\mathbf{X}})$. Let ${\mathbf{P}}\in\mathbb{S}^{n}$ denote the projection matrix onto the subspace $\textrm{Im}({\mathbf{X}})$ and let ${\mathbf{P}}_{\perp}\in\mathbb{S}^{n}$ denote the projection matrix onto the orthogonal complement of $\textrm{Im}({\mathbf{X}})$. Throughout the proof we let $\|{\cdot}\|_{1}$ and $\|{\cdot}\|_{2}$ denote the trace norm (sum of singular values) and the Euclidean (Frobenius) norm in some space of real (not necessarily symmetric) matrices, respectively. We begin by recalling two important facts. First, for any real (not necessarily symmetric) matrix ${\mathbf{A}}$ it holds that
$\|{{\mathbf{A}}}\|_{1}\leq\sqrt{\textrm{rank}({\mathbf{A}})}\|{{\mathbf{A}}}\|_{2}.$ (9)
Second, for any two real (not necessarily symmetric) matrices of appropriate dimension (such that the multiplication is defined) ${\mathbf{A}},{\mathbf{B}}$ it holds that
$\|{{\mathbf{A}}{\mathbf{B}}}\|_{2}\leq\|{{\mathbf{A}}}\|_{\infty}\|{{\mathbf{B}}}\|_{2},$ (10)
where we let $\|{{\mathbf{A}}}\|_{\infty}$ denote the largest singular value of ${\mathbf{A}}$.
Let us denote ${\mathbf{D}}={\mathbf{X}}-{\mathbf{Y}}$. It holds that,
$\|{{\mathbf{D}}}\|_{1}$ $=\|{({\mathbf{P}}+{\mathbf{P}}_{\perp}){\mathbf{D}}({\mathbf{P}}+{\mathbf{P}}_{\perp})}\|_{1}$
$\leq\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}+\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}+2\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}.$ (11)
We continue by upper-bounding each of the three terms in the RHS separately.
Note that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}})\leq k$. Thus, using Eq. (9) and then Eq. (10), and noting that $\|{{\mathbf{P}}}\|_{\infty}=1$, we have that
$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (12)
To bound the second term in the RHS of (B) we have,
$\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}$ $=\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}}\|_{1}\underset{(a)}{=}\textrm{Tr}({\mathbf{P}}_{\perp}{\mathbf{Y}})=\textrm{Tr}(({\mathbf{I}}-{\mathbf{P}}){\mathbf{Y}})$
$\underset{(b)}{\leq}\textrm{Tr}({\mathbf{X}})-\textrm{Tr}({\mathbf{P}}{\mathbf{Y}})=\textrm{Tr}({\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}})\leq\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}}}\|_{1}$
$\underset{(c)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2},$ (13)
where (a) holds since ${\mathbf{Y}}$ is positive semidefinite, and thus so is ${\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}$, (b) holds by the assumption that $\textrm{Tr}({\mathbf{Y}})=\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}=\textrm{Tr}({\mathbf{X}})$, and (c) is due to Eq. (12)
Finally, for the last term in the RHS of (B), noting that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp})\leq k$, using Eq. (9) and then Eq. (10) (and noting again that $\|{{\mathbf{P}}}\|_{\infty}=\|{{\mathbf{P}}_{\perp}}\|_{\infty}=1$), we have that
$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (14)
The lemma follows from plugging-in (12), (B), and (14) into the RHS of (B). ∎